OpenStudy (lostinlife15):

What is the equation of a line that has a slope of -4/5 that passes through (10, -3)?

OpenStudy (studygurl14):

y = mx + b m = slope = -4/5 Plug in the point (10,-3) for x and y, and solve for b

OpenStudy (sshayer):

y-y1=m(x-x1)